OpenStudy (anonymous):

Using complete sentences, explain each step in simplifying the ratio x^2+8x/x^2+10x+16 and provide the simplified form.

OpenStudy (anonymous):

someone please do this I have no clue

OpenStudy (anonymous):

Yeah me either I have been waiting forever.Ughhh

OpenStudy (anonymous):

I guess people always to the easy ones lol

OpenStudy (anonymous):

Yeahh nd then people like me wait like 4 days for an answer that I don't even need nymore because I gave up and guessed. lol

OpenStudy (whpalmer4):

Factor the numerator and the denominator. Cancel any common factors. Write down the answer :-)

OpenStudy (anonymous):

I literally just figured it out but there is another one I am have trouble with. Would you mind helping me?(:

OpenStudy (anonymous):

ohhhh...

OpenStudy (whpalmer4):

A similar problem, as an example of how to do this: \[\frac{3x^2-12x}{x^2+x-20} = \frac{3x(x-4)}{(x-4)(x+5)} = \frac{3x\cancel{(x-4)}}{\cancel{(x-4)}(x+5)} = \frac{3x}{x+5}\]

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

@shawna.marie sure, post the problem

OpenStudy (anonymous):

nice, how do you post these equations?

OpenStudy (anonymous):

Hold on.

OpenStudy (whpalmer4):

You can use the Equation button at the lower left, or typeset the material using LaTeX. To see what that looks like, select an equation by click-and-drag, then right-click on the selected material and choose Show Math As>TeX commands

OpenStudy (anonymous):

thnx :)

OpenStudy (whpalmer4):

@ynooran there's a LaTeX Practicing! subject that you can view on OpenStudy, too.

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

Are the equations 2x+4/3x+6 and 2/3 equivalent expressions? nd then it tells you to use complete sentences and all that other fun crap that no one cares about(:

OpenStudy (whpalmer4):

@ynooran upper left, just to the right of the OpenStudy logo, where it says "Mathematics" right now, click on that and scroll until you see it, then click again

OpenStudy (anonymous):

how to get latex?

OpenStudy (anonymous):

thnx bro g2g ill read your message later :)

OpenStudy (whpalmer4):

Well, you've got two fractions there, and you want to know if they are equivalent. The easy way to do that is to cross-multiply. Say we want to see if 2/3 and 4/8 are equivalent fractions: \[\frac{2}{3} =\frac{4}{8}\]\[2*8=3*4\]\[16=12\]That's not true, so the fractions are not equivalent. If we try that with your problem: \[\frac{2x+4}{3x+6} = \frac{2}{3}\]\[3*(2x+4) = 2*(3x+6)\]Is that a true statement for all values of \(x\)?

OpenStudy (whpalmer4):

@ynooran you just type the right characters to tell it you want to typeset something. "\""[" to start an equation, "\""]" to end one (without the quotation marks)

OpenStudy (whpalmer4):

@Shawna.marie I haven't left...

OpenStudy (anonymous):

Okay(:

OpenStudy (whpalmer4):

but I did just tell you how to do that problem — any questions?

OpenStudy (anonymous):

Ohh I didn't see that my bad.

OpenStudy (whpalmer4):

Report to the principal's office for your punishment :-)

OpenStudy (anonymous):

Lmao I do online schooling so I don't have a principal(: